OpenStudy (calculusxy):

How many two digit numbers are divisible by either 3 or 5?

OpenStudy (calculusxy):

@phi

OpenStudy (anonymous):

Didnt you ask this question? answers were given already

OpenStudy (kc_kennylau):

\[\frac{99-12}3+\frac{95-10}5-\frac{90-15}{15}\]

OpenStudy (calculusxy):

Yes!

OpenStudy (zpupster):

all of them

OpenStudy (kc_kennylau):

or (number of two digit numbers divisible by 3)+(number of two digit numbers divisible by 5)-(number of two digit numbers divisible by 15)

OpenStudy (phi):

This question has been answered a couple of times.

OpenStudy (anonymous):

common multiples of 3 and 5

OpenStudy (kc_kennylau):

and then of course, my expression is wrong xP

OpenStudy (kc_kennylau):

It should have been: \[\frac{99-12}3+1+\frac{95-10}5+1-\frac{90-15}{15}-1\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@kc_kennylau don't you mean +1 one the last one?

OpenStudy (kc_kennylau):

nope, coz it's in the parentheses :)

OpenStudy (anonymous):

O.O didn't catch that

OpenStudy (kc_kennylau):

xP